Find the slope of the graph of the relation 2x3 − 3xy + y3 = −1 at the point (2, −3).
negative eleven sevenths
5 over 7
negative 1 over 3
4 over 3

Answer by Alan3354(69443)

Find the slope of the graph of the relation 2x3 − 3xy + y3 = −1 at the point (2,−3).
----------------
Use ^ (Shift 6) for exponents.
---
2x^3 − 3xy + y^3 = −1
Differentiate implicitly.
6x^2dx - 3(ydx + xdy) + 3y^2dy = 0
2x^2dx - ydx - xdy + y^2dy = 0
2x^2dx - ydx = xdy - y^2dy
(2x^2 - y)dx = (x - y^2)dy
dy/dx = (2x^2 - y)/(x - y^2)
-------------
That's the slope, but (2,-3) is not a point on the curve.